# Hallucination to Truth: A Review of Fact-Checking and Factuality Evaluation in Large Language Models

### v1

@@ 6  Open issues and challenges
Table 5: Identified issues in LLM-based fact-checking and their implications.
@@
2. Advanced prompting (chain-of-thought, multi-agent) still leads to error cascades.
@@
Despite the advancement in fact-checking using LLMs, one common challenge remains the gap between model model-generated responses’ linguistic quality and factual accuracy. Today’s criteria of evaluation are inclined to appreciate language or text overlap with references, which can overlook basic fact errors. It guides to feedbacks that look questionable but sound good and inaccurate. These responses can receive arbitrarily high scores, which makes the system seem more accurate than it is [48, 58, 59].
Existing models often execute much better on small synthetic datasets, but often fail to generalize well to real-world circumstances. The reason lies in the limited datasets with low complexity, variation in topics, and multilingual texts. Consequently, the models do not handle the variability and nuance of real-world data across languages and topics, which implies the need for more realistic and broad training and test corpora [60, 80]. However, RAG techniques have a stronger evidence-based foundation in LLM, but retrieval is imperfect. Fact-correctness is often weakened by the quality of the retrieved information, noisy or irrelevant documents, and when there is insufficient context available. Similarly, advanced prompting techniques, such as chain-of-thought reasoning or multiagent collaborations, still must be precisely fine-tuned, but they also remain vulnerable to cascading errors in output generation [14, 77].
One of the areas with great potential that is still underexplored lies in the integration of LLMs with symbolic reasoning or structured logic-based systems. These systems can improve interpretability and fact resilience in fact-check pipelines. But work in this area is still in an early stage, and significant effort is needed to design scalable, proper architectures that can balance the respective strengths of neural and symbolic methods [76, 61]. Table 5 provides an organized summary of the existing issues and challenges discussed in this section.
@@ 7  Critical analysis of future research agendas
LLMs show great promise in automating fact checking, but their use also highlights a range of ongoing problems that still need attention. In the future, research must take a thoughtful and future-focused approach. If these models are to become trustworthy, accurate, and ethically sound tools in fact-checking, the gaps in today’s research, for example, the unreliability of retrieved evidence, the difficulty in verifying complex claims requiring multistep reasoning, and the challenge of mitigating factual hallucinations, need to be tackled head-on. What follows is a breakdown of key areas where further study could really make a difference, each pointing to where progress is most needed.
Table 6: Identified gaps from the proposed research questions (RQs), and future research paths in fact-checking with LLMs.
@@
Evaluation Framework Advancement. A fundamental imperative lies in transcending current evaluation metrics, which often inadequately capture the nuanced performance characteristics of LLMs in fact check tasks [3, 68]. Future research must prioritize the development of sophisticated, multi-dimensional frameworks. This includes establishing standardized metrics for explainability and interpretability that demonstrably correlate with human cognitive trust and facilitate diagnostic understanding of model failures [31, 38, 56]. Furthermore, the creation of dynamic evaluation suites (testing systems that can actively change and adapt over time, rather than relying on fixed, unchanging datasets) to rigorously test resilience against evolving misinformation tactics and sophisticated adversarial attacks is paramount [40], moving beyond static benchmarks. Currently, the development of precise metrics for fine-grained faithfulness and verifiable provenance tracking [50, 39] is crucial to ensure that LLM outputs are not merely plausible but demonstrably grounded in credible evidence.
@@
Logical Consistency, Reasoning, and Calibrated Trust. The ultimate efficacy of LLMs in fact checking is critically dependent upon their capacity for robust logical reasoning and their ability to engender warranted user trust. Future endeavors must explore formal verification methods (mathematically and logically ensuring that the model’s reasoning process is sound and its conclusions are consistent) to enhance the logical consistency of LLM outputs [17] and significantly deepen their reasoning capabilities for complex, multihop, and inferential claims [49]. The degree to which users believe the fact-checks offered by LLMs must also be thoroughly investigated. The goal of this study should be to appropriately "calibrate" or modify that trust to a suitable degree. The purpose is to encourage people to critically evaluate the information provided rather than relying too much on these automatic checks. [73, 56]
Multimodality and Multilinguality. Given that misinformation transcends unimodal (ie, English) texts, even though some studies have tried to incorporate multimodal and multilingual fact checking [43, 48, 32, 57], there is still a need for a significant expansion of LLM fact-checking capabilities. The development of robust multimodal systems capable of verifying claims that integrate textual, visual, and auditory information, and detecting sophisticated cross-modal manipulations is a key frontier [43, 31]. Equally vital are dedicated efforts to develop and evaluate information and effective and equitable fact-checking in a diverse spectrum of languages, with particular attention to resource-scarce linguistic contexts [32, 53].
@@ 8  Conclusion
The inclusion of LLMs in automated fact-checking is changing the field. These models are reshaping how we process and verify the overwhelming amount of information we face online. Our work lays out the current research in this space, drawing attention to five critical areas: how LLMs are evaluated, the problem of hallucinations (where models produce false information), the importance of data sources, various ways of improving performance, and the growing use of RAG. The findings highlight the fact that while this field is moving quickly and holds a lot of promise, there are serious issues that still need careful attention if we want these systems to work reliably.
One of the major achievements of this study is that it pulls together a broad snapshot of what is going on right now. It highlights a tricky balance: on the one hand, LLMs have the potential to improve the speed and quality of fact check. However, these systems are still limited, and fixing those limits is going to require a lot of effort from researchers. From the studies, one thing becomes clear: we need better tools to judge these systems. Since most existing tools only focus on whether a fact is right or wrong, there is now a growing need for tools that also look at how clearly the model explains things, whether it sticks to the logic, how well it shows where its answers came from, and how tough it is against tricky questions or misleading setups. From our review, it is evident that it is still difficult to ensure that LLMs always stick to real facts, especially when the situation is new or tricky. This improvement will require fundamental changes in the way these models are built and trained.
Although this review offers a comprehensive synthesis of the existing literature, its scope is inherently constrained by the rapid velocity of technological innovation within the LLM domain. Consequently, emerging preprint findings, advancements in proprietary models, and developing best practices may not be fully encapsulated. Furthermore, the predominant analytical lens has been applied to textual fact-checking, with multimodal and multilingual dimensions acknowledged as critical but less exhaustively explored frontiers. Looking ahead, future research must prioritize the development of more sophisticated, robust, and universally standardized evaluation benchmarks. Such benchmarks are urgently needed to assess not only the factual accuracy but also the logical coherence and soundness of LLM reasoning, the quality, utility, and persuasiveness of generated explanations, and the models’ resilience to a wide array of adversarial attacks and evolving misinformation tactics.

### v2

@@ 6  Open issues and challenges
Table 8: Identified issues in LLM-based fact-checking and their implications. This includes observed behaviors, underlying causes, and the significance of each issue for reliable deployment.
@@
2. Advanced prompting (CoT, multi-agent) still leads to error cascades.
@@
Despite the advancement in fact-checking using LLMs, one common challenge remains the gap between model-generated responses’ linguistic quality and factual accuracy. Today’s criteria of evaluation are inclined to appreciate language or text overlap with references, which can overlook basic fact errors. It guides to feedbacks that look questionable but sound good and is inaccurate. These responses can receive arbitrarily high scores, which makes the system seem more accurate than it is [48, 58, 59].
Existing models often execute much better on small synthetic datasets, but often fail to generalize well to real-world circumstances. The reason lies in the limited datasets with low complexity, variation in topics, and multilingual texts. Consequently, the models do not handle the variability and nuance of real-world data across languages and topics, which implies the need for more realistic and broad training and test corpora [60, 88]. Additionally, Siino et al. added that transformer-based models generally perform the worst in terms of standard deviation [84]. However, RAG techniques have a stronger evidence-based foundation in LLM, but retrieval is imperfect. Fact-correctness is often weakened by the quality of the retrieved information, noisy or irrelevant documents, and when there is insufficient context available. Similarly, advanced prompting techniques, such as CoT reasoning or multi-agent collaborations, still must be precisely fine-tuned, but they also remain vulnerable to cascading errors in output generation [14, 76, 119].
One of the areas with great potential that is still underexplored lies in the integration of LLMs with symbolic reasoning or structured logic-based systems. These systems can improve interpretability and fact resilience in fact-checking pipelines. But work in this area is still in an early stage, and significant effort is needed to design scalable, proper architectures that can balance the respective strengths of neural and symbolic methods [61, 75]. Table 8 provides an organized summary of the existing issues and challenges discussed in this section.
@@ 7  Critical analysis of future research agendas
LLMs show great promise in automating fact-checking, but their use also highlights a range of ongoing problems that still need attention. In the future, research must take a thoughtful and future-focused approach. If these models are to become trustworthy, accurate, and ethically sound tools in fact-checking, the gaps in today’s research, for example, the unreliability of retrieved evidence, the difficulty in verifying complex claims requiring multistep reasoning, and the challenge of mitigating factual hallucinations, need to be tackled head-on. What follows is a breakdown of key areas where further study could really make a difference, each pointing to where progress is most needed.
Table 9: Identified gaps from the proposed RQs, and future research paths in fact-checking with LLMs.
@@
Evaluation Framework Advancement. A fundamental imperative lies in transcending current evaluation metrics, which often inadequately capture the nuanced performance characteristics of LLMs in fact-checking tasks [3, 44]. Future research must prioritize the development of sophisticated, multi-dimensional frameworks. This includes establishing standardized metrics for explainability and interpretability that demonstrably correlate with human cognitive trust and facilitate diagnostic understanding of model failures [31, 38, 56, 120, 121, 122, 123, 124]. Furthermore, the creation of dynamic evaluation suites (testing systems that can actively change and adapt over time, rather than relying on fixed, unchanging datasets) to rigorously test resilience against evolving misinformation tactics and sophisticated adversarial attacks is paramount [40], moving beyond static benchmarks. Currently, the development of precise metrics for fine-grained faithfulness and verifiable provenance tracking [39, 50] is crucial to ensure that LLM outputs are not merely plausible but demonstrably grounded in credible evidence.
@@
Logical Consistency, Reasoning, and Calibrated Trust. The ultimate efficacy of LLMs in fact-checking is critically dependent upon their capacity for robust logical reasoning and their ability to engender warranted user trust. Future endeavors must explore formal verification methods (mathematically and logically ensuring that the model’s reasoning process is sound and its conclusions are consistent) to enhance the logical consistency of LLM outputs [17] and significantly deepen their reasoning capabilities for complex, multihop, and inferential claims [49]. The degree to which users believe the fact-checks offered by LLMs must also be thoroughly investigated. The goal of this study should be to appropriately "calibrate" or modify that trust to a suitable degree. The purpose is to encourage people to critically evaluate the information provided rather than relying too much on these automatic checks [56, 72]
Multimodality and Multilinguality. Given that misinformation transcends unimodal (ie, English) texts, even though some studies have tried to incorporate multimodal and multilingual fact-checking [32, 43, 48, 57], there is still a need for a significant expansion of LLM fact-checking capabilities. The development of robust multimodal systems capable of verifying claims that integrate textual, visual, and auditory information, and detecting sophisticated cross-modal manipulations is a key frontier [31, 43]. Equally vital are dedicated efforts to develop and evaluate information and effective and equitable fact-checking in a diverse spectrum of languages, with particular attention to resource-scarce linguistic contexts [32, 53].
@@ 8  Conclusion
The inclusion of LLMs in automated fact-checking is changing the field. These models are reshaping how we process and verify the overwhelming amount of information we face online. Our work lays out the current research in this space, drawing attention to five critical areas: how LLMs are evaluated, the problem of hallucinations (where models produce false information), the importance of data sources, various ways of improving performance, and the growing use of RAG. Furthermore, in addition to extensive textual fact-checking, our review also acknowledges assessment of multimodal and multilingual dimensions, recognized as core areas of contemporary research. The findings highlight the fact that while this field is moving quickly and holds considerable promise, there are significant challenges that still need careful attention if we want these systems to work reliably.
One of the major achievements of this study is that it pulls together a broad snapshot of what is going on right now. It highlights a tricky balance: on the one hand, LLMs have the potential to improve the speed and quality of fact-checking. However, these systems are still limited, and fixing those limits is going to require a lot of effort from researchers. It is essential to examine the practical utility of employing LLMs across all tasks, as smaller, task-specific models can often outperform them when properly optimized. From prior studies, one thing becomes clear: we need better tools to judge these systems. Since most existing tools only focus on whether a fact is right or wrong, there is now a growing need for tools that also look at how clearly the model explains things, whether it sticks to the logic, how well it shows where its answers came from, and how tough it is against tricky questions or misleading setups. From our review, it is evident that it is still difficult to ensure that LLMs always stick to real facts, especially when the situation is new or tricky. This improvement will require fundamental changes in the way these models are built and trained.
Although this review offers a comprehensive synthesis of the existing literature, its scope is inherently constrained by the rapid velocity of technological innovation within the LLM domain. Consequently, emerging preprint findings, advancements in proprietary models, and developing best practices may not be fully encapsulated. Looking ahead, future research must prioritize the development of more sophisticated, robust, and universally standardized evaluation benchmarks. Such benchmarks are urgently needed to assess the factual accuracy, logical coherence, soundness of LLM reasoning, quality, utility, and persuasiveness of generated explanations, as well as the models’ resilience to a wide array of adversarial attacks and evolving misinformation tactics.